how do i calculate the p-value on a calculator?Home Send to Gradebook < Prev Next > eTextbook SRIC Question 14 0.75/1 Show Attempt History Current Attempt in Progress Use t
Answer #1

Solution :

\mu = 4

\bar x = 4.8

s = 2.3

n = 15

This is the two tailed test .

Test statistic = t

= (\bar x - \mu ) / s / \sqrt n

= (4.8 - 4) / 2.3/ 15 \sqrt

= 1.347

P-value = 0.1993

\alpha =0.05

p = 0.1993 ≥ 0.05, it is concluded that the null hypothesis is not rejected.